# STAGNet: A Spatio-Temporal Graph and LSTM Framework for Accident Anticipation

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Accident prediction and timely warnings are vital for improving road safety and minimizing loss of lives and damage to property. Advanced Driver Assistance Systems (ADAS) play a crucial role in supporting human drivers, particularly in anticipating potential accidents before they happen. While many existing systems depend on a range of sensors such as LiDAR, radar, and GPS, relying solely on dashcam video input presents a more cost-effective and easily deployable solution. In this work, we tackle the challenging task of predicting imminent accidents using only dashcam footage. We analyze limitations in current methods and explore strategies to enhance their effectiveness. Our findings highlight significant potential to advance the current state-of-the-art in vision-based accident anticipation systems.
Analyzing dashcam video streams for extracting information for various driver assistance functions is a heavily researched area [1, 2]. In particular, several methods have been proposed to anticipate future accidents. Among them, graph neural network (GNN)-based methods have demonstrated superior performance [3, 4]. However, they do not utilize state-of-the-art techniques to extract spatio-temporal features from video frames containing fast-moving objects. Furthermore, they do not make optimal use of temporal attention. While some methods omit attention mechanisms [3, 5], others use spatial attention only [4]. In recent work, Takur et al. [6] apply temporal attention at the frame-level graph but not at the object-level graph. Furthermore, some of the video datasets widely used to train and test accident anticipation methods suffer from limitations such as video quality bias (e.g., CCD [5], A3D [7]) or a low proportion of ego-involved accidents (e.g., DAD [8][3]. However, evaluating these algorithms on unbiased datasets with a large proportion of ego-involved accidents is crucial, as such incidents are the ones an ADAS can directly mitigate rather than those solely involving other vehicles. Moreover, the generalization of models to datasets beyond those used to train them remains inadequately evaluated due to the use of biased datasets [2].
In this work, we attempt to overcome the aforementioned limitations of existing methods with the following contributions.
  • Using more advanced, spatio-temporal features (from SlowFast networks [9]) to represent the input video frames.
  • Aggregating frame-level features with a recurrent neural network.
  • Applying graph attention to the object-level graph.
@@
The architectural and feature extraction improvements we propose are simple, yet prove to be highly effective in practice.
@@ 2 Literature Review
Typically, accident anticipation methods rely on pretrained convolutional neural networks (CNNs) to detect objects in each video frame and extract features from the detected objects as well as the full frame. We will refer to these as object-level and frame-level features, respectively.
In one of the early works in accident anticipation, Chan et al. [8] proposed a method that feeds the extracted features through a spatial attention mechanism and a recurrent neural network (RNN). This enables the model to learn to focus on important objects based on cues such as their appearance, location, and motion. They have used VGG [10] to extract object-level features. For frame-level features, they have used VGG to extract features from each individual frame and IDT [11] to extract motion-related features from clips of 5 consecutive frames. Suzuki et al. [12] proposed a similar method with the main differences being the use of DeCAF [13] features instead of VGG and a quasi-recurrent neural network [14] for the RNN. Another approach similar to that in [8] was proposed by Karim et al. [15]. They use a similar attention mechanism to weight objects in the same frame (spatial attention). In addition, however, they use a temporal attention scheme where the last $M$ hidden states of the RNN are aggregated as a weighted sum. This allows the model to focus more on important frames in the video.
Several authors have proposed graph neural network-based accident anticipation models. In one of the pioneering works, Bao et al. [5] construct a graph for each video frame such that the nodes represent the objects detected in the frame and edges represent the distance between objects in pixel space. Initial node features are defined as a concatenation of object-level and frame-level features extracted using a pretrained VGG-16 model. The graph is then updated through a graph convolutional network (GCN) and a graph convolutional recurrent network (GCRN) [16] before going through a Bayesian neural network (BNN) [17] to produce the probability of an accident. The BNN also provides the uncertainty associated with the estimated probability. Mahmood et al. [3] proposed a similar graph neural network architecture but noticed that reliance on objects’ visual features may hinder the models’ ability to generalize to new scenes. To address this, they replaced object-level visual features with geometric features derived from bounding boxes and showed that it improves performance on datasets different from the one used to train the model (cross-dataset evaluation). Thakur et al. [6] proposed a nested graph approach, which further improved the performance. Their work forms the basis for our proposed work and we describe it in more detail in the next section.
In this section, we provided a brief review of the work most relevant to our work. We refer readers to [1, 2] for a more comprehensive review of the literature in accident anticipation.
@@ 3 Methodology
In this work, we propose an improved model to anticipate accidents from dashcam videos based on the Graph(Graph) framework [6]. Given a video consisting of a sequence of observed frames $(X_{1},X_{2},\dots,X_{N})$, the goal is twofold: (1) to predict whether an accident will occur, and (2) to detect it as early as possible. For each frame $X_{t}$, the model outputs a probability $p_{t}$ indicating the likelihood of an accident occurring in the future. For accident-positive cases, we define the Time-to-Accident (TTA) as $\tau=a-t$ for $t<a$, where $a$ is the start time of the accident and $t$ is the earliest frame such that $p_{t}>\alpha$, with $\alpha$ being a decision threshold.
The architecture of our model is illustrated in Fig. 1. Similar to the Graph(Graph) model [6], our model consists of three main modules. The first module is responsible for modeling the objects detected in video frames, while the second module extracts the global context of the video frames. The third module combines the information from the first two modules to provide the final prediction. In the rest of this section, we describe these modules, highlighting the changes we have made with respect to the Graph(Graph) model.
@@ 3.1 Spatio-Temporal Object Graph Learning
Not all objects captured by a dash-cam are equally important in assessing collision risk. Typically, objects that are closer to each other and/or closer to the ego-vehicle pose a higher risk of collision. Although it is challenging to accurately measure distances in pixel space with uncalibrated cameras, this module attempts to capture the relationships between traffic objects both within each frame (spatial) as well as across adjacent frames (temporal). The model learns to focus more on relevant objects through the use of spatio-temporal graph attention mechanisms.
We use YOLO [18] to detect objects and extract their bounding boxes from each video frame. For each bounding box, we extract its visual features using a pretrained VGG-16 [10] network, and the word embedding of the objects’ class label using GloVe [19]. For a given frame $X_{t}$, a spatio-temporal object graph $G^{\text{obj}}_{t}$ with $S$ nodes is constructed, where each node corresponds to a detected object and $S$ is the number of objects detected. The features of each node are initialized with a concatenation of the corresponding objects’ visual features $f^{\text{obj}}_{e}\in\mathbb{R}^{S\times d_{1}}$ and the word embedding of its class label $f^{\text{obj}}_{l}\in\mathbb{R}^{S\times d_{2}}$, where $d_{1}$ and $d_{2}$ are the visual feature and label embedding dimensions, respectively.
@@
where $s(i,j)$ is the cosine similarity between the object features of the $i$-th and $j$-th objects with labels $l_{i}$ and $l_{j}$, belonging to frames at time $t$ and tut-u, respectively. These edges ensure the flow of information between the objects of the same class across frames and avoid irrelevant information flow between objects from different classes.
The node embeddings are updated through Graph Attention Networks (GATv2) [20] using the following operations:
@@
where $\phi$ denotes a fully connected layer for dimensionality reduction, and $\cdot$] represents concatenation. Note that the original Graph(Graph) architecture used Graph Convolutional Networks (GCN) for this computation. However, we decided to use GATv2 instead as it allows each node in the graph to selectively attend to its neighbors based on its own feature representation and update it accordingly. This graph attention approach, particularly its improved version GATv2, has shown better performance than non-attentive methods such as GCNs on some of the benchmark graph datasets [20]. This attention mechanism can help our model learn to emphasize the most relevant object relationships in the sequence of scenes, improving prediction accuracy and supporting early accident anticipation.
3.2 Spatio-Temporal Global Feature Learning
While the spatio-temporal object graph captures fine-grained object interactions, extracting global contextual features from the entire frame sequence is crucial for providing a holistic understanding of the scene. Most previous approaches [5, 3, 15] rely on pretrained image models such as VGG16 [10] for global feature extraction. However, this does not capture temporal features that span across frames. Thakur et al. [6] extract global frame-level features with I3D [21], which derives features from a sequence of frames, thus capturing both spatial and temporal features. In this work, we utilize a pretrained SlowFast network [9] to extract frame-level spatio-temporal features. SlowFast networks have demonstrated better performance than I3D on video recognition benchmarks and have a dedicated path to process fast movements efficiently. This can be beneficial in processing dash-cam videos, where fast-moving objects are very likely.
To extract the global feature representation $f_{\text{fr}}^{t}$ for a frame at time $t$, we pass the current frame and the last-$u$ frames $X_{\text{seq}}=(X_{t-u},\dots,X_{t})$ into the SlowFast network, which produces a high-dimensional spatio-temporal representation. These features are then projected into a lower-dimensional space using a fully connected (FC) layer. To further enhance temporal modeling beyond what is captured by the SlowFast network, we pass the transformed features through a Long Short-Term Memory (LSTM) module:
@@ 3.3 Frame Graph Learning
To capture temporal relationships between video frames, a frame graph is constructed that leverages both global frame-level features and object-level features obtained from the previous stage. In this graph, each node corresponds to a frame at a particular time step $t$, and edges are directed from past frames to the current one, ensuring no future information is leaked backward in time. Each frame node is connected to its previous $k$ frames through $k$ directed edges. The adjacency matrix $A_{\text{fr}}$ for the frame graph is defined as:
$A^{\text{fr}}(i,j)=\begin{cases}1,&\text{if }i-j\leq k\\ 0,&\text{otherwise}\end{cases},$ (5)
@@
Figure 1: Architecture of the proposed STAGNet accident anticipation framework. Boxes with thicker borders represent the main components introduced in this work.
@@ 4 Experimental Evaluation
In this section, we detail the experimental setup, including the datasets used, the architecture of the network, the baseline methods for comparison, and the evaluation metrics. We then present and discuss the outcomes of the conducted experiments.
@@ 4.1.1 Datasets
We conduct our experiments on three benchmark dashcam video datasets: DAD, DoTA, and DADA. For the Dashcam Accident Dataset (DAD), we use the standard train-test split provided by the Graph(Graph) authors  [^1] https://github.com/thakurnupur/Graph-Graph, which is the same as that used in other works as well, ensuring a fair and consistent comparison with prior work. The DAD dataset comprises 1750 clips extracted from 678 dashcam videos, recorded at 20 frames per second, each lasting 5 seconds. Among these, 620 clips are labeled as accident (positive) and 1130 as normal (negative), with the accident occurring in the last 10 frames of each positive video. Note that this dataset primarily consists of non-ego-involved accidents (less than 10% ego-involved incidents) [15].
@@
Since the anticipation of ego-involved accidents is more important, we evaluate our model on two widely used accident datasets containing a large percentage of ego-involved accidents: DoTA [22] and DADA [23]. We extracted all the ego-involved accidents present in these datasets, which resulted in 745 positive and 745 negative videos from the DoTA dataset, and 965 positive and 1361 negative videos from the DADA dataset. All videos are of 5s duration (similar to DAD). A summary of the characteristics of these datasets is presented in Table 1 and more details can be found in [2]. To eliminate any bias from arbitrary dataset splitting, we adopt a $k$-fold cross-validation strategy, which allows for a more robust evaluation of model generalization. These datasets provide a diverse set of challenging driving scenarios essential for testing early accident anticipation performance in real-world conditions. To further evaluate model generalization, we conducted cross-dataset evaluation, where the model was trained on DoTA and tested on DADA, and vice-versa. Note that we down-sampled the videos in DADA from 30fps to 10fps for compatibility with those in DoTA to enable direct comparison of results.
@@
Note that we did not include the popular dataset CCD [5] in our evaluation due to its quality bias caused by the difference in video quality between accident and normal videos [3]. This bias can be exploited by models to classify videos based on quality rather than on accident cues. This is evident in the near-perfect performance of all models on CCD.
@@
DAD [8] 620 1130 1750 20
DoTA [22] 745 745 1490 10
DADA [23] 938 1361 2299 10

### v2

@@ Abstract
© 2025 IEEE. This is the author’s accepted manuscript of an article published in IEEE Access. The final published version is available at DOI: 10.1109/ACCESS.2025.3645127
@@ 1 Introduction
Accident prediction and timely warnings are vital for improving road safety by reducing injuries, fatalities, and property damage. Advanced driver assistance systems (ADAS) support human drivers by identifying potential hazards and are particularly valuable when they can anticipate accidents before they occur. Furthermore, accident anticipation provides an additional safety layer in autonomous vehicles [malawade2022spatiotemporal, fang2023vision, kumamoto2025aat]. While many existing systems depend on multiple sensors such as LiDAR, radar, and GPS, relying solely on dash-cam video offers a more cost-effective and easily deployable alternative. In this work, we address the problem of early accident anticipation using only dash-cam footage. We examine the limitations of existing methods and propose strategies to address them, thereby advancing vision-based accident anticipation.
Analyzing dash-cam video for driver assistance applications is a heavily researched area [fang2023vision, chitraranjan2025vision]. Among the many methods to anticipate future accidents, graph neural network (GNN)-based methods have demonstrated superior performance [thakur2024graph, mahmood2023new, malawade2022spatiotemporal]. In previous GNN-based methods [mahmood2023new, malawade2022spatiotemporal, bao2020uncertainty], graphs are built from individual video frames, and temporal dependencies are modeled later with some form of a recurrent neural network (RNN). In recent work, Takur et al. [thakur2024graph] embed temporal dependencies into the construction of the graph itself by adding edges from nodes that correspond to past frames. They also introduce graph attention to focus on important frames. However, their approach does not directly model sequential information. In this work, we introduce a Long Short-Term Memory (LSTM) network to explicitly capture the temporal dependencies in the sequence of frames and embed them into the graph’s nodes. The key distinction of our approach is the integration of explicit sequence modeling with spatio-temporal graph attention. Furthermore, the feature extractors used in existing methods either ignore temporal dependencies, as they operate solely at the individual frame level (VGG16 [simonyan2014very]), or are not tailored to videos with fast-moving objects (I3D [carreira2017quo]). We use SlowFast networks [Feichtenhofer2018SlowFastNF] to extract frame-level spatio-temporal features from sliding windows of the video stream. These networks capture both rapid motion dynamics (fast pathway) and slow-evolving spatial semantics (slow pathway). However, since SlowFast relies on temporal convolutions with small receptive fields (5 frames wide), it primarily captures short-term motion cues and is not designed to model long-term temporal dependencies. Introducing an LSTM module enables long-term temporal modeling. Similar ideas of augmenting temporal convolutions with LSTMs have produced improved performance in other domains [bi2021hybrid, hsu2022temporal].
Another critical problem lies in the data used to experimentally evaluate existing systems. Some video datasets widely used to train and test accident anticipation methods have video quality bias (e.g., CCD [bao2020uncertainty], A3D [yao2019unsupervised]) or a low proportion of ego-involved accidents (e.g., DAD [chan2017anticipating][mahmood2023new]. However, evaluating these algorithms on unbiased datasets with a large proportion of ego-involved accidents is crucial, as such incidents are the ones that an ADAS can directly mitigate, rather than those solely involving other vehicles. Moreover, the generalization of models to datasets beyond those used to train them remains inadequately evaluated due to the use of biased datasets [chitraranjan2025vision].
In summary, our contributions aimed at addressing the aforementioned problems are as follows.
  • Encoding the input video frames with advanced spatio-temporal features capable of capturing rapid motion.
  • Combining sequence modeling with graph attention by employing an LSTM that aggregates spatio-temporal features across frames before passing them on to the GNN.
  • Applying graph attention to the object-level graphs.
  • Providing a reduced version of the model for faster inference.
@@
These architectural and feature extraction choices are simple, yet prove highly effective in practice and enable our model to outperform the state-of-the-art. We also show that the reduced version of our network is considerably faster than existing methods while still achieving higher average precision and time-to-accident values in predicting ego-involved accidents.
@@ 2 Literature Review
Typically, accident anticipation methods rely on pretrained convolutional neural networks (CNNs) to detect objects in each video frame and extract features from the detected objects and the full frame. We will refer to these as object-level and frame-level features, respectively.
In one of the early works, Chan et al. [chan2017anticipating] proposed a method that feeds the extracted features through a spatial attention mechanism and a recurrent neural network (RNN). This allows the model to emphasize important objects based on cues such as appearance, location, and motion. They used VGG16 [simonyan2014very] to extract object-level features. For frame-level features, they used VGG16 to extract features from each individual frame and IDT [wang2013action] to extract motion-related features from clips of 5 consecutive frames. Suzuki et al. [suzuki2018anticipating] proposed a similar method, with the main differences being the use of DeCAF [donahue2014decaf] features instead of VGG and a quasi-recurrent neural network [bradbury2016quasi] for the RNN. Another approach similar to [chan2017anticipating] was proposed by Karim et al. [karim2022dynamic]. They use a similar attention mechanism to weight objects in the same frame (spatial attention). In addition, however, they use a temporal attention scheme where the last $M$ hidden states of the RNN are aggregated as a weighted sum. This allows the model to focus more on important frames in the video.
Several authors have proposed graph neural network-based accident anticipation models. In one of the pioneering works, Bao et al. [bao2020uncertainty] construct a graph for each video frame such that the nodes represent the objects detected in the frame, and the edges represent the distances between objects in pixel space. Initial node features are defined as a concatenation of object-level and frame-level features extracted using a pretrained VGG16 model. The graph is then updated through a graph convolutional network (GCN) and a graph convolutional recurrent network (GCRN) [seo2018structured] before passing through a Bayesian neural network (BNN) [neal2012bayesian] that produces the probability of an accident and its associated uncertainty. Mahmood et al. [mahmood2023new] used a similar graph neural network architecture but observed that reliance on visual object features may reduce cross-dataset generalization and proposed replacing them with geometric features derived from bounding boxes, resulting in improved cross-dataset performance. Thakur et al. [thakur2024graph] proposed a nested graph approach that further improved performance. Their work forms the basis for our proposed method, and we describe it in more detail in the next section.
The above studies summarize the methods most relevant to our approach. We refer readers to [fang2023vision, chitraranjan2025vision] for a more comprehensive review of accident anticipation literature.
@@ 3 Methodology
In this work, we propose an improved model to anticipate accidents from dashcam videos based on the Graph(Graph) framework [thakur2024graph]. Given a sequence of observed video frames $(X_{1},\dots,X_{t-1},X_{t})$, the goal is twofold: (1) to predict whether an accident will occur or not, and (2) to detect it as early as possible. For each frame $X_{t}$, the model outputs a probability $p_{t}$ indicating the likelihood of an accident occurring in the future. For accident-positive cases, the Time-to-Accident (TTA) is defined as $\tau=a-t$ for $t<a$, where $a$ is the start time of the accident and $t$ is the earliest frame such that $p_{t}>\alpha$, with $\alpha$ being a decision threshold.
The architecture of our model is illustrated in Figure 1. Similar to the Graph(Graph) model [thakur2024graph], our model consists of three main modules. Module A captures the spatio-temporal relationships between detected objects, while module B extracts the global context of the video. Module C combines information from the other two modules to provide the final prediction. In the rest of this section, we describe these modules, highlighting the changes we have made with respect to the Graph(Graph) model.
@@ 3.1 Spatio-Temporal Object Graph Learning
Not all objects captured by a dash-cam contribute equally to collision risk. Objects that are closer to each other and/or to the ego-vehicle pose a higher risk of collision. Although precise distance estimation is difficult with uncalibrated cameras, this module attempts to capture the relationships between traffic objects, both within each frame (spatial) and across adjacent frames (temporal). The model learns to focus more on relevant objects through the use of spatio-temporal graph attention mechanisms.
We use YOLO [wang2024yolov10realtimeendtoendobject] to detect objects and extract their bounding boxes from each video frame. For each bounding box, we extract visual features using a pretrained VGG16 [simonyan2014very] network and the word embedding of the objects’ class label using GloVe [pennington2014glove]. For a given frame $X_{t}$, a spatio-temporal object graph $G^{\text{obj}}_{t}$ with $S$ nodes is constructed, where each node corresponds to a detected object, and $S$ is the number of objects detected. The features of each node are initialized with a concatenation of the corresponding objects’ visual features $f^{\text{obj}}_{e}\in\mathbb{R}^{S\times d_{1}}$ and the word embedding of its class label $f^{\text{obj}}_{l}\in\mathbb{R}^{S\times d_{2}}$, where $d_{1}$ and $d_{2}$ are the visual feature and label embedding dimensions, respectively.
@@
where $s(i,j)$ is the cosine similarity between the object features of the $i$-th and $j$-th objects with labels $l_{i}$ and $l_{j}$, belonging to frames at time $t$ and tut-u, respectively. These edges propagate information between objects of the same class across frames and avoid irrelevant information flow between objects from different classes.
The node embeddings are updated through Graph Attention Networks (GATv2) [brody2021attentive] using the following operations:
@@
where $\phi$ denotes a fully connected layer for dimensionality reduction, and $\cdot$] represents concatenation. Note that the original Graph(Graph) architecture used Graph Convolutional Networks (GCN) for this computation. We adopt GATv2 instead of GCNs because its attention mechanism allows each node to attend selectively to the most relevant neighbors. This graph attention approach, particularly its improved version GATv2, has shown better performance than non-attentive methods such as GCNs on some of the benchmark graph datasets [brody2021attentive]. This attention mechanism can help our model learn to emphasize the most relevant object relationships in the sequence of scenes, improving prediction accuracy and supporting early accident anticipation.
3.2 Spatio-Temporal Frame-level Feature Learning
While the spatio-temporal object graph captures fine-grained object interactions, extracting global contextual features from the entire frame sequence is crucial for providing a holistic understanding of the scene. Most previous approaches [bao2020uncertainty, mahmood2023new, karim2022dynamic] rely on pretrained image models such as VGG16 [simonyan2014very] for global feature extraction. However, this does not capture temporal features that span across frames. Thakur et al. [thakur2024graph] extract global frame-level features with I3D [carreira2017quo], which derives features from a sequence of frames, thus capturing both spatial and temporal features. In this work, we utilize a pretrained SlowFast network [Feichtenhofer2018SlowFastNF] to extract frame-level spatio-temporal features. SlowFast networks have demonstrated better performance than I3D on video recognition benchmarks and have a dedicated path to process fast movements efficiently. This is useful in processing dash-cam videos, which often contain fast-moving objects.
To extract the global feature representation $f_{\text{fr}}^{t}$ for a frame at time $t$, we pass the current frame and the last-$u$ frames $X_{\text{seq}}=(X_{t-u},\dots,X_{t})$ through the SlowFast network, which produces a high-dimensional spatio-temporal representation. These features are then projected into a lower-dimensional space using a fully connected (FC) layer. To further enhance temporal modeling, we pass the transformed features through a Long Short-Term Memory (LSTM) module:
@@ 3.3 Frame Graph Learning
To capture temporal relationships between video frames, a frame graph is constructed that leverages both global frame-level features and object-level features obtained from the previous stage. In this graph, each node corresponds to a frame at a particular time step $t$, and edges are directed from past frames to the current one, ensuring that no future information is leaked backward in time. Each frame node is connected to its previous $k$ frames through $k$ directed edges. The adjacency matrix $A_{\text{fr}}$ for the frame graph is defined as:
$A_{\text{fr}}(i,j)=\begin{cases}1,&\text{if }i-j\leq k\\ 0,&\text{otherwise}\end{cases},$ (5)
@@
Figure 1: Architecture of the proposed STAGNet accident anticipation framework. Boxes with thicker borders represent the main components introduced in this work. The components within the red dashed rectangle are pruned in the light-weight model, STAGNet-Lite.
3.4 Model Pruning
We note that extracting object-level features from each bounding box and constructing object graphs is computationally expensive and may be unnecessary for ego-involved accident anticipation. Since the ego-vehicle is not detected as an object, its relationship to other objects is not modeled in the object graphs. Hence, we hypothesized that the spatio-temporal frame-level feature learning module and its corresponding frame-level graph are sufficient to predict ego-involved accidents. Therefore, we propose a lightweight version of our model (STAGNet-Lite) by pruning the spatio-temporal object graph learning module and its corresponding frame-level graph for the efficient anticipation of ego-involved accidents. The red dashed rectangle in Figure 1 encapsulates the pruned components.
@@ 4 Experimental Evaluation
In this section, we first present the details of the experimental setup, datasets used, the baseline methods, and the evaluation metrics. We then discuss the outcomes of the experiments conducted.
@@ 4.1.1 Datasets
We conducted experiments on three benchmark dash-cam video datasets: DAD [chan2017anticipating], DoTA [yao2022dota], and DADA [fang2019dada].
@@
The Dashcam Accident Dataset (DAD) consists of 1,750 five-second clips derived from 678 dashcam videos recorded at 20 fps. Among them, 620 clips are annotated as ‘accident’ (positive) and 1,130 as ‘normal’ (negative). In all positive clips, the accident event occurs within the final 10 frames. Note that this dataset primarily consists of non-ego-involved accidents (less than 10% ego-involved incidents) [karim2022dynamic]. We used the standard train-test split provided by the authors of Graph(Graph)  [^1] https://github.com/thakurnupur/Graph-Graph, which is the same as that used in other work as well. This ensures a fair and consistent comparison with prior work.
Since the anticipation of ego-involved accidents is more important, we evaluated our model on two widely used accident datasets containing a large percentage of ego-involved accidents: DoTA [yao2022dota] and DADA [fang2019dada].
From the DADA dataset, we selected all clips labeled as ego-involved accidents and extracted five-second segments such that the accident onset occurred within the last two seconds, with the exact start time randomized (as in CCD). This produced 938 positive samples. Negative samples were obtained by extracting five-second segments from normal driving periods preceding the accident windows. Although we considered each video in the dataset as a potential source for a negative sample, those lacking a sufficiently long pre-accident segment were excluded. This procedure yielded 1361 negative samples.
Positive and negative clips from the DoTA dataset were extracted following a similar procedure. Based on the duration of the video clips and the accident onset times, we were able to extract 745 negative samples depicting normal driving scenes. To construct a balanced dataset [^2] A balanced dataset was used to prevent inflated AP scores that can occur when positive samples greatly outnumber negative ones [saito2015precision]., we randomly selected an equal number of videos containing ego-involved accidents to produce 745 positive samples. Note that Mahmood et al. [mahmood2023new] have followed a similar approach to select videos from DoTA and reported similar numbers of videos. The decision to set the duration of each extracted clip to five seconds was based on other popular datasets, e.g., DAD and CCD.
We adopt a $k$-fold cross-validation strategy for DoTA and DADA, which provides a more robust evaluation of model performance. During cross-validation, we group clips so that any samples from the same source video are in the same fold.
A summary of the characteristics of these datasets is presented in Table 1, and more details can be found in [chitraranjan2025vision]. The videos and the relevant scripts can be found at this link.
@@
These datasets provide a diverse set of challenging driving scenarios that are essential for testing early accident anticipation performance in real-world conditions. To further evaluate model generalization, we conducted a cross-dataset evaluation, where the model was trained on DoTA and tested on DADA, and vise versa. Note that we down-sampled the videos in DADA from 30 fps to 10 fps for compatibility with those in DoTA to enable direct comparison of results.
@@
Note that we did not include the popular dataset CCD [bao2020uncertainty] in our evaluation due to its quality bias, caused by the difference in video quality between accident and normal videos [mahmood2023new]. This bias can be exploited by models to classify videos based on quality rather than on accident cues. This is evident in the near-perfect performance of all models on CCD.
@@
DAD [chan2017anticipating] 620 1130 1750 20
DoTA [yao2022dota] 745 745 1490 10
DADA [fang2019dada] 938 1361 2299 10